# Mitigating “Epistemic Debt” in Generative AI-Scaffolded Novice Programming using Metacognitive Scripts

### v1

@@ 5.4. Industry Implications: Realization of Epistemic Debt
The risks identified in our laboratory setting find a direct parallel in recent industry events, validating the Epistemic Debt hypothesis at scale. The Clawdbot/Moltbot incident of early 2026 (Reuters, 2026) serves as a primary case study of the dangers of high utility coupled with low corrective competence. In that event, thousands of developers deployed personal AI agents using generated configurations they did not fully understand.
@@
5.5. Fragile Velocity vs. Sustainable Velocity
A common critique of metacognitive interventions is that they impede development speed. However, our data suggests this is a false dichotomy. While the Explanation Gate introduced a time penalty of approximately 14 minutes in Group C, this cost was largely subsidized by the generation speed of the LLM.
We frame this trade-off as a shift from Fragile Velocity to Sustainable Velocity. The VibeCheck intervention ensures that the velocity gained through AI is maintainable by the human developer. By enforcing the “Teach-Back” protocol, the tool ensures that the intrinsic load of the task is offloaded rather than outsourced.
@@ 5.6. Self-Scaffolding and Interactional Stances
The outlier analysis in RQ3 provides a final, critical insight: accumulation of Epistemic Debt is not solely a function of the tool, but of the user’s interactional stance. Successful participants in the Unrestricted group naturally adopted a “Consultant Stance.” This suggests that while advanced learners may possess the metacognitive discipline to treat AI as a peer, novices require structural friction to prevent them from falling into the “Contractor Stance.”
Future learning systems must, therefore, move beyond passive code generation toward active metacognitive mediation. By automating the verification of student explanations through an LLM-as-a-Judge framework (Choma et al., 2025), we can scale such interventions to meet the demands of learners in the current GenAI era.
@@ 6.2. External Validity
Our study focused exclusively on novice programmers (students and bootcampers). It is possible that for expert developers, who already possess robust internal mental models, the Explanation Gate would represent purely extraneous load with no germane benefit. The “Epistemic Debt” we observed is likely a function of the novice-AI dyad, and results may not generalize to experienced programmers who utilize AI for boilerplate tasks rather than core logic. In either case, approaching the problem from first principles, i.e., defining germane, intrinsic and extrinsic loads will allow for the delineation of offloading and outsourcing for the respective populations, and designing appropriate scaffolds, as a result. We posit that our study, grounded in cognitive load theory, while not externally face valid nevertheless provides directional support for the design of scaffolds for experienced programmers as well.
@@ 6.3. Construct Validity
We used “Repair Success” as a proxy for Corrective Competence. Critics might argue that this measures debugging skill rather than understanding. However, as noted in Section 3.4.2, because the bug was a regression of the user’s own code integrated moments prior, we believe it remains a valid measure of cognitive ownership.
@@ 7. Conclusion
This study has demonstrated that the democratization of software creation via GenAI carries the latent risk of an industry-wide collapse of competence. However, this breakdown is not inevitable. By deliberately introducing Metacognitive Friction into the development loop, we can leverage the immense productivity gains of LLMs while ensuring that the humans at the helm remain capable of maintaining the complex systems they build. To prevent the mass production of unmaintainable software, we must ensure that the price of velocity is not the loss of comprehension.
@@ 8.1. Adaptive Scaffolding and Instructional Fading
The current implementation of the VibeCheck Explanation Gate utilizes a static threshold (SOLO Level 3). Future work should explore adaptive friction, where the complexity of the required explanation “fades” as the system detects increasing learner competence. Drawing on the Expertise Reversal Effect (Kalyuga, 2009), a system that transitionally moves from high-friction scaffolds to frictionless utility could prevent the “Fragile Expert” syndrome while eventually granting the developer the high-velocity benefits of an unrestricted environment.
@@ 8.2. Longitudinal Learning and Mental Model Decay
Our experiment measured corrective competence immediately following a construction task. However, the long-term persistence of schemas formed under metacognitive friction remains unknown. Future longitudinal studies should investigate whether “buying back” epistemic debt through teach-back protocols leads to higher retention of architectural principles over months of professional practice, or if the “Consultant Stance” naturally erodes as users face the time-pressures of industrial production.
@@ 8.3. Collaborative Human-AI-Human Systems
While this study focused on the programmer-AI dyad, modern software engineering is a social activity. Research is needed to determine how epistemic debt impacts code reviews and collaborative debugging. If a senior developer reviews code authored by a novice “Vibe Coder,” does the lack of cognitive ownership in the junior dev create a bottleneck in the team’s collective corrective competence? Extending the Explanation Gate to include collaborative “Teach-Forward” protocols between humans could mitigate these social bottlenecks.
@@ 8.4. Equity, Diversity, and the AI Experience Gap
A critical area for future inquiry is the impact of Generative AI on underrepresented groups in Computer Science. Currently, Black and Hispanic students earn close to 20% of CS degrees in the United States, combined, (Siek and Batten, 2025; National Center for Education Statistics, 2023) yet they often report higher rates of “imposter syndrome” and lower access to early-career mentorship. Generative AI has the potential to act as a “leveling agent”; however, if unrestricted AI usage leads to a disproportionate accumulation of epistemic debt among learners with less existing social capital, it may inadvertently widen the “competence gap” during technical interviews. Future research must quantify whether scaffolded AI tools like VibeCheck provide a more equitable pathway to senior-level architectural understanding for students from diverse socio-economic and racial backgrounds.

### v2

@@ 5.4. Industry Implications: Realization of Epistemic Debt
The risks identified in our laboratory setting find a suggestive parallel in recent industry events, lending ecological plausibility to the epistemic debt hypothesis. The Moltbook incident of early 2026 (Reuters, 2026) serves as a primary case study of the dangers of high utility coupled with low corrective competence. In that event, thousands of developers deployed personal AI agents using generated configurations they did not fully understand.
@@
5.5. The Speed-Comprehension Trade-off
A common critique of metacognitive interventions is that they impede development speed. However, our results suggest this is a false dichotomy. This critique also rests on a productivity assumption that may itself be flawed: Shen and Tamkin (Shen and Tamkin, 2026) find in a professional programming context that AI-assisted developers showed no statistically significant improvement in task completion time, yet suffered a 17% reduction in subsequent skill assessment scores ($d=0.74$). The implication is that unrestricted AI may not even deliver the speed gain it is presumed to trade off against learning. While the Explanation Gate introduced a time penalty of approximately 14 minutes in Group C, this cost was largely subsidized by the generation speed of the LLM.
Crucially, the speed gained through AI assistance in Group C was matched by the comprehension required to maintain it. The VibeCheck intervention distinguishes between velocity that is fragile (fast but unmaintainable) and velocity that is sustainable (fast and cognitively owned). By enforcing the teach-back protocol, the tool ensures that the intrinsic load of the task is offloaded rather than outsourced.
@@ 5.6. Self-Scaffolding and Interactional Stances
The outlier analysis in RQ3 provides a final, critical insight: accumulation of epistemic debt is not solely a function of the tool, but of the user’s interactional stance. Successful participants in the Unrestricted group naturally adopted a consultant stance. This suggests that while advanced learners may possess the metacognitive discipline to treat AI as a peer, novices require structural friction to prevent them from falling into the contractor stance. Concordant evidence from industry usage data is provided by Anthropic’s Economic Index (Anthropic, 2026), which independently classifies real-world interactions into automation (directive, full-delegation) and augmentation (feedback loop, task iteration, validation, learning) categories. Their analysis finds that high-tenure users gravitate away from directive automation toward iterative augmentation, precisely the shift from contractor to consultant stance that our intervention enforces structurally for novices.
The contractor stance observed in Group B corresponds to the emotional dependence on AI interlocutors identified by Yan et al. (Yan et al., 2026). These participants did not merely use the tool; they offloaded their critical agency to it. Our Explanation Gate broke this dependence by reintroducing the metacognitive friction necessary to restore the learner’s epistemic agency. A complementary taxonomy of stances emerges independently in Shen and Tamkin (Shen and Tamkin, 2026), who identify six interaction patterns in professional developers learning a new API with AI assistance. Their three low-scoring patterns (AI Delegation, Progressive AI Reliance, and Iterative AI Debugging) correspond directly to the contractor stance, while their high-scoring patterns (Generation-Then-Comprehension, Hybrid Code-Explanation, and Conceptual Inquiry) map onto the consultant. Most tellingly, their Conceptual Inquiry pattern (asking only explanatory questions, never requesting generated code) achieved quiz scores of 65–86% while matching the completion speed of full delegation, providing direct independent evidence that interactional stance, not tool access, is the determinant of learning outcomes.
Future learning systems must, therefore, move beyond passive code generation toward active metacognitive mediation. By automating the verification of student explanations through an LLM-as-a-Judge framework (Choma et al., 2025), we can scale such interventions to meet the demands of learners in the current GenAI era. This imperative extends beyond competence. Lodge and Loble (Lodge and Loble, 2026) argue that education is not merely a cognitive undertaking but a process of formation: novices become thinking practitioners by struggling genuinely with hard problems. Our findings suggest that the Explanation Gate does not only prevent competence collapse; it preserves the conditions under which that formation can occur.
@@ 6.2. External Validity
Our study focused exclusively on novice programmers (students and bootcampers). It is possible that for expert developers, who already possess robust internal mental models, the Explanation Gate would represent purely extraneous load with no germane benefit. The epistemic debt we observed is likely a function of the novice-AI dyad, and results may not generalize to experienced programmers who utilize AI for boilerplate tasks rather than core logic. That said, Dell’Acqua et al. (Dell’Acqua et al., 2026) demonstrate that even highly skilled, incentivized knowledge workers (elite BCG consultants) show significant over-reliance failure on outside-frontier tasks, suggesting that the underlying mechanism is not confined to novices. The implication is that scaffolding interventions of the type we study may generalize further up the skill spectrum than our participant sample implies. In either case, a CLT-grounded approach, defining intrinsic, extraneous, and germane loads for each population, can guide the design of appropriate scaffolds. While our study was conducted with a specific novice population, it provides directional support for scaffolding design applicable to experienced programmers as well.
@@ 6.3. Construct Validity
We used repair success as a proxy for Corrective Competence. Critics might argue that this measures debugging skill rather than understanding. However, as noted in Section 3.4.2, because the bug was a regression of the user’s own code integrated moments prior, we believe it remains a valid measure of cognitive ownership.
6.4. Ecological Validity
The experiment used a single, controlled task (a React.js Course Scheduler) completed within a two-hour session. Real-world epistemic debt accumulates across months of production work, across diverse technology stacks, and in collaborative codebases. Our study measures an acute form of debt incurred within one session; whether this proxy generalizes to chronic, career-long debt accumulation is an open question. We believe the controlled setting is a strength for causal inference, but future work should replicate the core finding in longitudinal field deployments.
@@ 7. Conclusion
This study has demonstrated that the democratization of software creation via GenAI carries the latent risk of an industry-wide collapse of competence. However, this breakdown is not inevitable. By deliberately introducing metacognitive friction into the development loop, we can leverage the immense productivity gains of LLMs while ensuring that the humans at the helm remain capable of maintaining the complex systems they build. In our experiment, the Explanation Gate reduced epistemic debt by 41.7 points (from 69.3 to 27.6) at a velocity cost of only 14.2 minutes, a cost largely subsidized by the LLM’s generation speed. To prevent the mass production of unmaintainable software, we must ensure that the price of velocity is not the loss of comprehension. The VibeCheck plugin and the full experimental protocol are openly available as a replication artifact to support this line of inquiry.
@@ 8.1. Adaptive Scaffolding and Instructional Fading
The current implementation of the VibeCheck Explanation Gate utilizes a static threshold (SOLO Level 3). Future work should explore adaptive friction, where the complexity of the required explanation fades as the system detects increasing learner competence. Drawing on the “expertise reversal effect” (Kalyuga, 2009), a system that transitionally moves from high-friction scaffolds to frictionless utility could prevent the fragile expert syndrome while eventually granting the developer the high-velocity benefits of an unrestricted environment.
@@ 8.2. Longitudinal Learning and Mental Model Decay
Our experiment measured corrective competence immediately following a construction task. However, the long-term persistence of schemas formed under metacognitive friction remains unknown. Future longitudinal studies should investigate whether buying back epistemic debt through teach-back protocols leads to higher retention of architectural principles over months of professional practice, or if the consultant stance naturally erodes as users face the time-pressures of industrial production.
@@ 8.3. Collaborative Human-AI-Human Systems
While this study focused on the programmer-AI dyad, modern software engineering is a social activity. Research is needed to determine how epistemic debt impacts code reviews and collaborative debugging. If a senior developer reviews code authored by a novice vibe coder, does the lack of cognitive ownership in the junior dev create a bottleneck in the team’s collective corrective competence? Extending the Explanation Gate to include collaborative teach-forward protocols between humans could mitigate these social bottlenecks.
@@ 8.4. Equity, Diversity, and the AI Experience Gap
A critical area for future inquiry is the impact of Generative AI on underrepresented groups in Computer Science. Currently, Black and Hispanic students combined earn close to 20% of CS degrees in the United States (Siek and Batten, 2025; National Center for Education Statistics, 2023), yet they often report higher rates of imposter syndrome and lower access to early-career mentorship. Lodge and Loble (Lodge and Loble, 2026) identify a related metacognitive equity gap: the cognitive risks of unstructured AI fall disproportionately on novices who lack the domain knowledge and self-regulation skills needed to use AI beneficially, compounding existing disadvantage. Generative AI has the potential to act as a leveling agent; however, if unrestricted AI usage leads to a disproportionate accumulation of epistemic debt among learners with less existing social capital, it may inadvertently widen the competence gap during technical interviews. Future research must quantify whether scaffolded AI tools like VibeCheck provide a more equitable pathway to senior-level architectural understanding for students from diverse socio-economic and racial backgrounds.
Artifact Availability
The VibeCheck VS Code/Cursor extension, the Course Scheduler experimental task suite (starter scaffold, reference implementation, 12-assertion Puppeteer grading harness, and logic bomb version), and the complete three-condition study replication protocol are openly available at https://github.com/sreecharansankaranarayanan/vibecheck